carmelle:

Six times a number decreased by 10 is eight. Find the number… A. 2 B. 3 C. 4 D. 5

madiurmom:

so we can solve that 6x-10=8

XioGonz:

6x - 10 = 8 6x = 18 x = 3

madiurmom:

we move ten to the right side so it would be 6x-10=8 +10 +10 then -10 +10 = 0 and 8+10=18 so 6x=18 then divide by lcd

xxDeppressionxx:

6•3=18-10=8

carmelle:

ty @XioGonz and @madiurmom

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle ty @XioGonz and @madiurmom \(\color{#0cbb34}{\text{End of Quote}}\) Np

madiurmom:

your welcome : )

xxDeppressionxx:

Np